determine the transformations of the function f(x)=−4x+1 from the parent function.


1
Expert's answer
2022-02-24T11:02:03-0500

The parent function is "g(x)=x."

Stretch the graph of "y=g(x)" vertically by a factor of "4"


"h(x)=4x"

Reflect the graph of "y=h(x)" in the x-axis


"u(x)=-4x"

Shift the graph of "y=u(x)" upward "1" unit


"f(x)=-4x+1"